Three tennis balls are tightly packed in a can.(circular cylinder). What fraction of the volume of the can is occupied by the tennis balls?
Answer by stanbon(75887)

Three tennis balls are tightly packed in a can.(circular cylinder). What fraction of the volume of the can is occupied by the tennis balls?
-----------
Volume of each ball is (4/3)pi*r^3
Total ball volume = 3(4/3)pi*r^3 = (4*pi*r^3)
------------------------
Volume of the can:
Area of the base = pi*r^2
--
Height of the can = 6r
---
Volume of the can 6r*pir^2 = 6pir^3
----------------------------------
Desired Fraction: 6pir^3/4pir^3 = 3/2
=====================================
Cheers,
Stan H.